OpenStudy (anonymous):

is this anwser 4???????? The y-intercept of the line whose equation is 2x + y = 4 is -2 2 4

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok thank u i was wounding if i did it right....

OpenStudy (anonymous):

welcome

OpenStudy (anonymous):

how did u get that anwser??????????

OpenStudy (anonymous):

help witch of the answer is right??????????/5 or -2

OpenStudy (lifeisadangerousgame):

Why would it be -2??

OpenStudy (rina.r):

wait let me do it on my note book.

OpenStudy (texaschic101):

put you equation in y = mx + b form and the b position is the y intercept

OpenStudy (rina.r):

b coz of y=mx+c

OpenStudy (rina.r):

yeah put it in this equation

OpenStudy (texaschic101):

2x + y = 4 y = -2x + 4 In y = mx + b, the b position is the y intercept. Therefore, the 4 is the y intercept. Understand ?

OpenStudy (anonymous):

i would do it this way: 2x+y=4 2(0)+y=4 y=4 the answer is 4 and i think tht the right answer..

OpenStudy (rina.r):

yeah its 4

OpenStudy (texaschic101):

true..that is another way. To find y intercept, sub in 0 for x. To find x intercept, sub in 0 for y.

OpenStudy (anonymous):

thank u guys i did the way khadeeja did.....